Please help me in this problem, I need this tommorow: How much pure alcohol must be added to 25 mL of water to produce 50% alcoholic solution?
Found 2 solutions by stanbon, MathLover1:
Answer by stanbon(75887)

How much pure alcohol must be added to 25 mL of water to produce 50% alcoholic solution?
----
Equation:
alcohol + alcohol = alcohol
---
x + 0*25 = 0.5(x+25)
---
Multiply thru by 100
100x + 0 = 50x+50*25
50x = 50*25
x = 25mL (amt. of pure alcohol needed)
==========================
Cheers,
Stan H.

Answer by MathLover1(20850)


25+ml of pure alcohol must be added to 25 mL of water to produce 50% alcoholic solution